Grade 9Mechanics

A ball is thrown vertically upwards with an initial velocity such that it can reach a maximum height of 15m.If ,at the same instance ,a stone is dropped from a height of 15m,find the ratio of distance travelled by them when they cross each other.

Razz kumar
6 Years agoGrade 9

1 Answer

Vikas TU
ApprovedApproved Tutor Answer6 Years ago
Dear student 
velocity of projection of ball:v1=√2gh
=√2×9.8×15
=√294
=17.15 m/s
Time taken to cross each other;
Distance travelled by ball + stone = h
v1t−1/2gt^2+1/2gt^2=h
t=h/v1
=15/17.15 
=0.875 s
Ratio of distances travelled:
Ratio=v1t−1/2gt^2 / 12gt^2
=17.15×0.875−1/2×9.8×0.875^2 / 0.5×9.8×0.875^2=15−3.75/3.75
=11.25 / 3.75
=3:1